# Rydberg Atomic Quantum Receivers for Classical Wireless Communication and Sensing

### v1

@@ II-C1 RAQR Noise Sources
There are several factors that contribute to the noise of RAQRs. The photon shot noise and the electrical (Johnson-Nyquist) noise from the photodetector may increase the noise floor, making it difficult to detect weak signals. The laser noise plays a crucial role in determining the sensitivity of RAQR. The frequency noise of lasers can be minimized by frequency stabilization techniques, where the laser is locked to the mode of a high-finesse ultra-low expansion (ULE) cavity. The intensity noise of lasers can be reduced similarly by implementing a servo feedback loop and a balanced photodetection scheme. The vibrational instability of mirrors may contribute to the mechanical noise. All these can be mitigated by innovative engineering techniques.
II-C2 Transit Time and Doppler Broadening
The atoms in the vapor cell only interact with laser light during the so called transit time. After this they leave the interaction region due to their thermal motion (kinetic energy) at room temperature. This leads to the broadening of spectral lines termed as transit time broadening.
@@ II-C2 Transit Time and Doppler Broadening
An atom moving in (against) the direction of light beam will experience an energy level shift i.e., redshifted (blueshifted) with respect to the light frequency. When this frequency is resonant to an atomic transition, light gets absorbed by the atoms obeying the same (Gaussian) velocity distribution of the atoms. These different velocities result in the broadening of spectral lines, termed as Doppler broadening. In RAQR, there is a mismatch in wavelengths for the counter-propagating probe and coupling laser beams which accounts for the Doppler-broadened Rydberg-EIT peak (shown in Fig 3(c)). The broadening is proportional to the residual Doppler mismatch and it is a few MHz.
II-C3 Quantum Projection Noise [3]
It is also known as the quantum shot noise and often referred to as Standard Quantum Limit (SQL), since it sets a lower limit to the RF signal strength that the the RARQ can receive. This noise arises from the quantum fluctuations of the measurement process. The SQL scales as $\propto 1/\sqrt{NT_{r}T_{i}}$, where $N$ is the number of uncorrelated Rydberg atoms participating in RF reception, $T_{i}$ is the integration time and $T_{r}$ is the coherence time of the RAQR. The typical SQL sensitivity is in the range of a few hundred pV/cm/$\sqrt{\text{Hz}}$.
In order to break through the SQL, one has to harness quantum-enhanced measurement systems, such as the complex quantum states of light (Schrodinger cat state, squeezed state, etc) and interacting many-body states. A sensitivity enhancement has been shown recently for a many-body, non-equilibrium Rydberg atomic system [10].
II-C4 Collisional Effects [3]
@@
The interaction or collision of Rydberg atoms with the vapor-cell wall becomes significant at roughly micron-scale distances, since there are boundary conditions imposed on the atomic dipoles, as they approach the wall. There are other collisions in this system, like Rydberg atom - Rydberg atom collisions and ground-state - Rydberg atom collisions, that play an important role. All of these collisions cause dephasing of the quantum states and reduce the coherence time $T_{r}$ of the RAQR receiver. It can be shown that when all these dephasing effects are taken into account, the coherence time $T_{r}$ and hence the sensitivity reaches the maximum value for Rydberg states in the range of $n=45,\cdots,60$.
@@ II-C5 Noise from RF Background
RAQRs cannot differentiate between an RF target signal and the background electromagnetic noise. The change in the response of RAQR in the presence of bandwidth-limited white Gaussian RF noise has been theoretically modelled and verified experimentally. Additionally, by combining it with machine learning models, the RAQR can be trained to receive a particular RF signal of given bandwidth even if the output spectroscopic signal is heavily distorted by noise.
@@ II-D Comparison to Conventional RF Receivers
II-D1 Different Operating Principle
Conventional RF receivers employ antennas for capturing the incident EM waves and to generate currents or voltages. By contrast, the RAQRs realize RF reception via Rydberg excitation created by driving laser beams in an atom vapor cell.
II-D2 SI-Traceable Receiver With Self-Calibration
@@
II-D3 Direct Passband to Baseband Conversion
Current RF receivers require tightly-integrated RF front-end components (e.g., mixers, amplifiers, and filters, etc.) for down-converting the passband RF signal to baseband signals, requiring complex integrated circuits, especially for large-scale systems. By contrast, the measurement scheme of RAQRs allow direct down-conversion of RF signals, bypassing the complex integrated circuitry, thus facilitating a simple, compact structure.
II-D4 Different Receiver Impairments
Traditional RF receivers suffer from Gaussian thermal noise and hardware impairments, while RAQRs suffer from a multitude of different impairments - ranging from the noise of lasers and photodetectors to noise arising from broadening and dephasing effects in atomic systems, as explained in Section II-C. The sensitivity for RAQRs is set by the quantum shot noise limit (SQL), whereas for conventional RF receivers the sensitivity is determined by the thermal noise limit. Ref. [11] has argued that the quantum shot noise will duly drop below the thermal noise floor of the conventional RF receivers, when the symbol rate is lower than 1-2 Msymbol/s, provided that all other experimental noise sources are absent.
@@ III RAQR Enabled Quantum-Aided Wireless Communication and Sensing
III-A Verified Capabilities
  • C2: Phase Detection. It can also retrieve the phase of an RF signal using the superheterodyne principle of Section II-B2.
  • C3: Polarization Detection. The polarization of an RF signal can be detected from the spectroscopic EIT signal, which depends on the angular difference between the laser polarization and the RF wave vector [12]. Moreover, an atomic system relying on polarization selectivity becomes feasible with the aid of metallic structures.
  • C4: Modulation Detection. The amplitude and phase detection capabilities of RAQRs allow the detection of various amplitude/frequency modulation (AM/FM) schemes [13, 12], phase-shift keying (PSK) (BPSK and QPSK [14, 8]), and of quadrature amplitude modulation (QAM) [14].
  • C5: Multiband/Continuous-Band Detection. Simultaneously detecting RF signals in multiple bands becomes feasible using a single RAQR. This can be realized by using different Rydberg final states, where each final state is coupled with a distinct RF band. Alternatively, populating a vapor cell with different atoms and matching each atom species with different laser wavelengths can sense different RF bands [13]. The multiple bands can span over a continuous frequency range through adjacent Rydberg resonance tuning [8].
  • C6: Spatial Displacement and Direction Detection. The spatial displacement of a moving target or angle-of-arrival (AoA) of an incident RF signal[7]) can be detected by RAQRs. By recording the phase-change of the RF signal caused by target movements, one can deduce spatial displacement [15]. and by acquiring the phase-difference of the RF signal between two locations, the AoA can be estimated [7].
@@ III-A Verified Capabilities
                                                              Theoretical studies from the physics community
   Ref.
EIT model
Sensitivity
Atomic
motions
Atomic
collisions
Number of atoms
Vapor cell influence
             Main Contributions
[3]
$\times$
$\surd$
$\times$
$\surd$
$\surd$
$\surd$
(dielectric constant)
Using a bright resonance prepared within an EIT window, it is possible to achieve
high sensitivities
$\sim$
1 µV/cm/$\sqrt{\text{Hz}}$ and detect small RF electric fields
1 µV/cm.
                                                              Experimental studies from the physics community
@@
Laser power
(probe/coupling)
Laser wavelength
(probe/coupling)
@@
Standard
@@
[12]
Standard
C1, C3, C4 (AM,
FM, pulsed RF)
F1, F2,
F4
133 Cs
6S 1/2 →6P 3/2
nD 5/2 →(n+1)P 3/2
N/A
852 nm / 510 nm
12.605 GHz,
2.5 GHz
A portable & commercial Rydberg atomic
RF electric-field measurement instrument
is first reported.
[14]
Superheterodyne
C1, C2, C4
(BPSK, QPSK,
QAM)
F1, F3
133 Cs
6S 1/2 →6P 3/2
34D 5/2 →35P 3/2
41.2 µW /
48.7 mW
850.53 nm /
511.148 nm
LO: 19.629 GHz,
IF: 500 kHz
& 1 MHz
The BPSK, QPSK, and QAM modulated
signals are able to be received by RAQRs.
The instantaneous bandwidth of Rydberg
atoms is about 1 MHz - 10 MHz.
[6]
Superheterodyne
C1, C2
F1
133 Cs
6S 1/2 →6P 3/2
47D 5/2 →48P 3/2
120 µW / 34 mW
852 nm / 510 nm
LO: 6.94 GHz,
IF: 150 kHz
Experimentally illustrate a sensitivity of
55 nV/cm/$\sqrt{\text{Hz}}$ and show the minimum
detectable field of 780 pV/cm.
[7]
Superheterodyne
C1, C2, C6 (AoA)
F1, F5
133 Cs
6S 1/2 →6P 3/2
59D 5/2 →59P 3/2
96 µW / 60 mW
852.35 nm /
509.26 nm
19.18 GHz
Superheterodyne Rydberg atomic sensor
is experimentally verified to detect AoA.
[8]
Superheterodyne
C1, C2, C4
(QPSK), C5
(continuous-band)
F1, F2,
F5
87 Rb
5S 1/2 →5P 3/2
70D 5/2 →71P 3/2
→71S 1/2
3.8 µW / 60.4 mW
780 nm / 480 nm
100 MHz -
1.4 GHz
Detect over 1 GHz continuous-band with
best sensitivity of 1.5 µV/cm/$\sqrt{\text{Hz}}$ and
80 dB linear dynamic range.
                                                              Applications to wireless communication and sensing from the communication community
@@ III-B Distinctive Features
  • F1: Extremely-High Sensitivity. Compared to conventional RF receivers, RAQRs exhibit ultra-high sensitivity in detecting RF signals [3, 6], as discussed in Sections II-B1 and II-B2.
  • F2: Broadband Tunability. RAQRs allow a multitude of options in choosing different atomic energy levels combined with different electron transitions to realize the detection of RF signals across a broad frequency range using a single receiver [12, 2].
  • F3: Narrowband Selectivity. The atomic response has a narrowband selectivity, where a narrow instantaneous bandwidth ($\leq 10$ MHz) around the carrier frequency can be realized for coupling between adjacent Rydberg states [14]. This facilitates EM interference-free receptions.
  • F4: Ultra-Wide Input Power Range. RAQRs are capable of receiving not only very weak RF signals, but also very strong RF signals up to several kV/m [12], thereby expanding the linear dynamic range. This ensures the receiver’s operation in high-intensity RF environments.
  • F5: Ultra-High Scalability. RAQRs exhibit an ultra-high scalability and an ability to form advanced receivers (e.g., multiband or continuous-band receivers and arrays) using a single vapor cell. A multiband/continuous-band receiver architecture is facilitated by methods stated in C5 [13, 8]. Construction of a receive array is feasible via splitting laser beams into multiple parallel rays and arranging for each probe-coupling beam pair to propagate through a unique location of the cross-section of a vapor cell [7]. Additionally, the vapor cell can be integrated with metallic structures or embedded into antennas to form a multi-functional receiver (e.g., amplitude-phase-polarization detection) [12].
@@ III-C State-of-the-Art in RAQRs
Next, we review several representative works (see TABLE I), spanning from theoretical aspects to prototype experiments, to applications in communication and sensing, from both the physics and communication communities.
III-C1 Theoretical Studies From the Physics Community [3]
The theoretical studies on RAQRs are mainly from the physics community. A salient aspect of these studies is establishing the Rydberg-EIT signal model by considering a specific 4-level ladder scheme. Another aspect is unveiling the fundamental limitations in approaching the SQL and exploring the optimum sensitivity. Last but not least, several practical limitations must also be considered, such as the noise from lasers and photodetectors, broadening, dephasinng and collisional effects, the influences of vapor cell geometry, etc.
III-C2 Experimental Studies From the Physics Community [14, 6, 13, 12, 7, 8]
@@
Another prevalent branch in the physics community is experimentally designing novel atomic structures to support new capabilities, demonstrate specific features, explore new sensitivity limits, and so forth. The advances of RAQRs have confirmed the feasibility of realizing powerful RF receivers, which will pave the way for more complex upper-layer communication and sensing applications.
III-C3 Applications to Wireless Communication and Sensing From the Communication Community [15]
Attracted by the appealing advantages and inspired by recent advances of RAQRs, upper-layer applications in wireless communication and sensing started to emerge. The authors of [15] demonstrated its superiority in sensing moving objects.
Figure 4: Superheterodyne RAQR based wireless communication transceiver.
IV Rydberg Atomic Quantum SISO and MIMO Schemes

### v2

@@ II-C1 RAQR Noise Sources
Several factors contribute to the noise of RAQRs. Among them, the photon shot noise and the electrical noise from the photodetector (PD) may increase the noise floor. The laser noise plays a crucial role in determining the sensitivity of RAQR. The frequency noise of lasers can be minimized by frequency stabilization techniques, where the laser is locked to the mode of a high-finesse ultra-low expansion cavity. The intensity noise of lasers can be reduced similarly by implementing a servo feedback loop and a balanced photodetection scheme. The vibrational instability of mirrors may contribute to the mechanical noise. All these can be mitigated by innovative engineering techniques.
II-C2 Transit Time and Doppler Broadening [2]
@@
The atoms in the vapor cell only interact with laser beams during the so called transit time. They will leave the interaction region due to their thermal motion (kinetic energy) at room temperature. This leads to the broadening of spectral lines termed as transit time broadening. An atom moving in (against) the direction of light beam will experience an energy level shift i.e., redshifted (blueshifted) with respect to the light frequency. When this frequency is resonant to an atomic transition, light gets absorbed by the atoms obeying the same (Gaussian) velocity distribution of the atoms. These different velocities result in the broadening of spectral lines, termed as Doppler broadening. In RAQRs, there is a mismatch in wavelengths for the counter-propagating probe and coupling laser beams which accounts for the Doppler-broadened Rydberg-EIT peak. The broadening is proportional to the residual Doppler mismatch.
II-C3 Collisional Effects [2]
@@
The interaction or collision of Rydberg atoms with the vapor-cell wall becomes significant at roughly micron-scale distances, since there are boundary conditions imposed on the atomic dipoles, as they approach the wall. There are other collisions in this system, like Rydberg atom - Rydberg atom collisions and ground-state - Rydberg atom collisions, that play an important role. All of these collisions cause dephasing of the quantum states and reduce the coherence time $T_{r}$ of the RAQR receiver. It has been shown in [2] that when all these dephasing effects are taken into account, the coherence time $T_{r}$ and hence the sensitivity reaches the maximum value for Rydberg states in the range of $n=45,\cdots,60$.
II-C4 Quantum Projection Noise [5]
It is also known as the quantum shot noise and often referred to as SQL, since it sets a lower limit to RAQRs’ sensitivity. This noise arises from the quantum fluctuations of the measurement process. The SQL scales as $\propto 1/\sqrt{NT_{r}T_{i}}$, where $N$ is the number of uncorrelated Rydberg atoms participating in RF reception, $T_{i}$ is the integration time and $T_{r}$ is the coherence time of the RAQR. The typical SQL sensitivity is on the order of pV/cm/$\sqrt{\text{Hz}}$ [8].
@@ II-C5 Noise from RF Background
RAQRs cannot differentiate between an RF target signal and the background electromagnetic noise. The change in the response of RAQR in the presence of bandwidth-limited white Gaussian RF noise has been theoretically modelled and verified experimentally, as seen from [3] and reference [98] of [3].
@@ II-D Comparison to Conventional RF Receivers
II-D1 Operating Principle
Traditional RF receivers capture incident electric fields by using antennas and generate the corresponding currents. The size of antennas requires to match the wavelength of the detected electric field. By contrast, RAQRs realize RF reception via Rydberg atoms created by driving laser beams in a vapor cell. The size of RAQRs is independent to the wavelength.
II-D2 SI-Traceability and Calibration
@@
II-D3 Passband-Baseband Conversion
Classical RF receivers, exploiting the popular homodyne/heterodyne/superheterodyne RF structures, require tightly-integrated RF front-end components (e.g., mixers, amplifiers, filters) for down-converting the passband RF signal to baseband, requiring complex integrated circuits, especially for large-scale systems [1]. By contrast, the measurement scheme of RAQRs allow direct down-conversion of RF signals [15], bypassing the complex integrated circuitry, thus facilitating a simple, compact structure.
II-D4 Receiver Impairments
Classical antenna-based receivers suffer from Gaussian thermal noise and electronic hardware impairments, while RAQRs suffer from several different impairments, ranging from the noise of lasers and PDs to noise arising from broadening and dephasing effects in atomic systems, as explained in Section II-C. The sensitivity of RAQRs is set by the SQL [2, 3, 4, 5], whereas the sensitivity of their counterparts is determined by the thermal noise limit, as discussed in [2] and reference [39] of [2].
II-D5 Angular Reception Ranges
Antenna (array-) based receivers can only receive impinging RF signals from a restricted spatial angle, e.g., half-plane range to the greatest extent for antenna arrays. By contrast, RAQRs are capable of receiving signals from wider spatial angles with lesser restrictions, facilitating the reduction of sensor deployments.
II-D6 Mutual Coupling Effects
Conventional antenna arrays typically face negative mutual coupling effects, which become more serious as antenna arrays become dense and large. The mutual coupling effects may cause considerable challenges for receiver designs. Fortunately, RAQRs are less susceptible to mutual coupling as Rydberg sensors do not re-emit the impinging RF signal, facilitating simplified receiver designs.
@@ III RAQR Enabled Quantum-Aided Wireless Communication and Sensing
                                                              Theoretical and experimental studies from the physics community
@@
Laser (probe/coupling) Power        Wavelength
@@
Superheterodyne
C1, C2, C4
(BPSK, QPSK, QAM)
F1, F3
133 Cs
6S 1/2 →6P 3/2
34D 5/2 →35P 3/2
41.2 µW /
48.7 mW
850.53 nm /
511.148 nm
LO: 19.629 GHz,
IF: 500 kHz & 1 MHz
The BPSK, QPSK, and QAM modulated
signals are able to be received by RAQRs.
[8]
Superheterodyne
C1, C2
F1
133 Cs
6S 1/2 →6P 3/2
47D 5/2 →48P 3/2
120 µW /
34 mW
852 nm /
510 nm
LO: 6.94 GHz,
IF: 150 kHz
Experimentally illustrate a sensitivity of
55 nV/cm/$\sqrt{\text{Hz}}$ by superheterodyne.
[9]
OLA,
superheterodyne
C1, C2
F1
133 Cs
6S 1/2 →6P 3/2
44D 5/2 →45P 3/2
30 µW /
135 mW
852 nm /
509 nm
LO: 8.57 GHz
Experimentally illustrate a sensitivity of
19 nV/cm/$\sqrt{\text{Hz}}$ by proposing the OLA.
[10]
PAS
C1, C2, C3
F1, F6
85 Rb
5S 1/2 →5P 3/2
60D 5/2 →59F 7/2
N/A
780 nm /
480 nm
LOx 10.6633 GHz,
LOy 10.66327 GHz,
LOz 10.6628 GHz
Experimentally illustrate a sensitivity of
57 µV/m/$\sqrt{\text{Hz}}$ and 66 µV/m/$\sqrt{\text{Hz}}$ for
the horizontal and vertical polarizations.
[11]
ARRT
C1, C5
(continuous-band)
F1, F2,
F4, F6
133 Cs
6S 1/2 →6P 3/2
68S 1/2 →67P 3/2 (68P 3/2 )
→65D 5/2 (66D 5/2 )
N/A
850 nm /
509 nm
11.25 GHz -
13.9 GHz
The ARRT scheme was demonstrated for
detecting continuous-band signals.
[12]
ARRT,
superheterodyne
C1, C2, C4
(QPSK), C5
(continuous-band)
F1, F2,
F4, F6
87 Rb
5S 1/2 →5P 3/2
70D 5/2 →71P 3/2
→71S 1/2
3.8 µW /
60.4 mW
780 nm /
480 nm
100 MHz -
1.4 GHz
Detect over 1 GHz continuous-band with
best sensitivity of 1.5 µV/cm/$\sqrt{\text{Hz}}$ and
80 dB linear dynamic range.
[13]
CRS
C1, C5
F1, F2,
F6
133 Cs
6S 1/2 →6P 3/2 →19D 5/2
17F 7/2 →17G 9/2 →17H 11/2
→17I 13/2 →17K 15/2
17L 17/2 →17M 19/2
20 µW /
21 mW
852 nm /
519 nm
0.607 THz,
34.9 GHz,
6.01 GHz, 1.85 GHz,
700, 310, 128 MHz
The CRS scheme was shown for simul-
taneously detecting multiband signals.
[14]
DRFS,
superheterodyne
C1, C2, C4
(QPSK), C5
F1, F2,
F6
87 Rb
5S 1/2 →5P 3/2
56D 5/2 →57P 3/2
(54F 7/2 , 59P 3/2 , 52F 7/2 )
3.8 µW /
60.4 mW
780 nm /
480 nm
1.72, 12.11 GHz,
27.42, 65.11 GHz,
115.75 GHz
The DRFS scheme was shown for simul-
taneously detecting five RF tones from
1.7 GHz to 116 GHz.
[15]
MAS
@@ 34.9 GHz,
                                                              Wireless communication and sensing applications from the communication community
@@ 28 GHz
[18]
Superheterodyne
C1, C2, C4
F1, F2
133 Cs
6S 1/2 →6P 3/2 →47D 5/2
→48P 3/2 (45D 5/2 →46P 3/2 ),
(43D 5/2 →44P 3/2 ), (40D 5/2
→41P 3/2 ), (66S 1/2 →66P 3/2 ),
(63S 1/2 →63P 3/2 )
29.8 µW
(9.4 µW)
/ 17 mW
852 nm /
510 nm
6.9458, 7.9752,
9.2186, 11.6187,
13.4078, 15.5513 GHz
Propose a RAQR-SISO scheme, signal
models, and provide performance analysis
for wireless communication and sensing.
[19]
Superheterodyne
C1, C2, C6 (DOA)
F1, F2
133 Cs
6S 1/2 →6P 3/2
47D 5/2 →48P 3/2
29.8 µW
/ 17 mW
852 nm /
510 nm
6.9458 GHz
Propose a RAQ-MIMO wireless sensing
scheme, signal model, and a multi-target
DOA estimation algorithm.
III-A Verified Capabilities
  • C2: Phase Detection. It can also retrieve the phase of an RF signal by harnessing the superheterodyne principle of Section II-B2.
  • C3: Polarization Detection. The polarization of an RF signal can be detected from the spectroscopic EIT signal by exploiting the PSS scheme of Section II-B4.
  • C4: Modulation Identification. The amplitude and phase extraction capabilities allow RAQRs to identify RF signal waveforms relying on different modulations, such as analog modulation schemes: amplitude/frequency modulation (AM/FM) [15], and digital modulation schemes: phase shift keying (PSK) (BPSK, QPSK [7, 12]) as well as quadrature amplitude modulation (QAM) [7].
  • C5: Multiband/Continuous-Band Detection. Simultaneously detecting multiband RF signals becomes feasible using a single RAQR. This can be realized by the CRS [13], DRFS [14], and MAS [15] schemes of Section II-B6. The multiple bands can span over a continuous frequency range through the ARRT scheme of Section II-B5.
  • C6: Spatial Displacement and Direction Detection. The spatial displacement and direction-of-arrival (DOA) of a moving target can be detected by RAQRs. By recording the RF signal phase-change caused by target movements, one can deduce spatial displacement [17]. By acquiring the phase-differences of the RF signal between different locations, DOA estimation can be realized [19].
@@ III-B Distinctive Features
  • F1: Extremely-High Sensitivity. Compared to traditional RF receivers, RAQRs exhibit ultra-high sensitivity in detecting RF signals, as discussed in Section II-B1, II-B2, and II-B3. The sensitivity can be on the order of $<$ nV/cm/$\sqrt{\text{Hz}}$ [3, 9], surpassing the 1.5 nV/cm/$\sqrt{\text{Hz}}$ sensitivity of metal antennas [3].
  • F2: Broadband Tunability. RAQRs allow a multitude of options in choosing different atomic energy levels combined with different electron transitions to realize the detection of RF signals across a broad frequency range (DC to THz) using a single receiver [5].
  • F3: Narrowband Selectivity. For a standard RAQR scheme, the atomic response has a narrowband selectivity ($\leq$ 10 MHz [5]), where only a narrow instantaneous bandwidth around the carrier frequency can be realized for coupling between adjacent Rydberg states [2, 5].
  • F4: Enhanced Instantaneous Bandwidth. The instantaneous bandwidth can be significantly extended by exploiting the ARRT scheme [11], which can be up to $100$ MHz [3] and reference [39] of [3], as discussed in II-B5. This facilitates RAQRs in receiving RF signals having a large continuous bandwidth.
  • F5: Ultra-Wide Input Power Range. RAQRs are capable of receiving not only very weak signals, but also very strong signals up to several kV/m [5]. This ensures the receiver’s operation in high-intensity RF environments.
  • F6: Ultra-High Scalability. RAQRs exhibit an ultra-high scalability to form advanced receivers (e.g., arrays, multiband, continuous-band, and PSS receivers) using a single vapor cell based on the schemes discussed in Section II-B. These functionalities can be potentially integrated with a single vapor cell to realize multi-functional receivers.
Figure 4: Superheterodyne RAQR based wireless reception scheme, signal model, and simulation results.
@@ III-C State-of-the-Art in RAQRs
Next, we review several representative works from both the physics and communication communities (seen in TABLE I).
III-C1 Theoretical and Experimental Studies From the Physics Community
@@
The studies on RAQRs from the physics community can be classified into two branches. The first branch is from a theoretical perspective, aiming to analyse the fundamental performance limit of SQL and practical impairments on RAQRs [2, 5], and to construct specific quantum response models for novel RAQR concepts, such as many-body Rydberg atom [3]. Another branch is experimentally designing novel atomic structures to support new capabilities, demonstrate specific features, explore new sensitivity limits, and so forth [8, 9, 10, 11, 12, 13, 14, 15, 7]. The advances of RAQRs have confirmed the feasibility of realizing powerful RF receivers, which paves the way for more complex upper-layer communication and sensing applications.
III-C2 Wireless Communication and Sensing Applications From the Communication Community
Inspired by recent advances of RAQRs, wireless communication and sensing applications started to emerge. Reference [17] experimentally demonstrated RAQRs’ superiority in sensing moving objects. Then, the work [18] took the lead in constructing useful signal models and analysing the superior performance of RAQR-aided wireless systems. Furthermore, the work [19] exploited a Rydberg receiver array for realizing multi-target DOA estimation by proposing a corresponding DOA estimation algorithm.
IV RAQ-SISO and RAQ-MIMO Schemes